# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v1

@@ I Introduction
Our formula is valid for any continuous symmetry described by a compact Lie group. To illustrate its wide applicability, we provide four examples. First, since the asymptotic conversion theory among pure states has been completed in prior research exclusively for the case of $U(1)$ [21], we confirm that this result immediately follows from our formula. Second, we provide a necessary and sufficient condition for conversion to be reversible. As a consequence, we affirmatively prove the Marvian-Spekkens conjecture, which had been unsolved for the past decade until today. Third, we analyze examples for $U(1)$ and $SU(2)$ in which the asymptotic conversion is irreversible. The conversion rate for the example in $SU(2)$ is shown to be consistent with the prior study on a specific class of pure states in [2]. Fourth, we calculate the conversion rate for finite groups. Although a finite group is not a Lie group, it is shown that our formula is applicable by appending a trivial $U(1)$ phase. We find a divergence in conversion rate, reproducing the prior result on the approximate conversion rate in RTA for finite groups in [34]. Based on our proof of the direct part, we also provide an intuitive explanation of the reason why the approximate conversion rate can diverge in RTA for finite groups.
@@
The rest of this paper is organized as follows: We first review the framework of RTA and the definition of the QGT in Sec. II. The main result of this paper, i.e., the approximate asymptotic conversion rate among pure states in RTA for an arbitrary Lie group, is presented in Sec. III. In Sec. IV, we show that the QGT is asymmetry monotone for pure states in one-shot and asymptotic conversion. As a consequence, we prove an inequality providing a bound for the asymptotic conversion rate, i.e., the converse part of the main theorem. In Sec. V, we show the optimality of this bound, i.e., the direct part of the main theorem. In Sec. VI, we apply our formula for RTA for several setups, providing a unified understanding of prior studies on conversion rates in RTA for different groups [2, 4, 6, 37, 21, 34], and proves the Marvian-Spekkens conjecture [6, 4].
II Setup and notations
II.1 The resource theory of asymmetry (RTA)
@@
II.2 Quantum geometric tensor (QGT)
@@
for $\mu=1,\cdots,\dim G$, where this corresponds to $L(A_{\mu})$ with the Lie algebra representation $L(A)\coloneqq-\mathrm{i}\left.\frac{\mathrm{d}}{\mathrm{d}t}U(e^{\mathrm{i}tA})\right|_{t=0}$. Then the QGT for |$\ket{\xi(\bm{\lambda})}$ at |$\ket{\psi}$, which we simply denote by $\mathcal{Q}^{\psi}$, is equal to the non-symmetrized covariance matrix for $\bm{X}\coloneqq\{X_{\mu}\}_{\mu=1}^{\dim G}$:
@@ IV.1 Monotonicity of QGT in one-shot conversion
$\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}=\frac{1}{1-q}V(\psi,O)+\frac{1}{q}V(\psi,O)^{*},$ (12)
@@
In literature, the proof of the monotonicity of monotone metric is sometimes provided under the assumption that $\rho$ is invertible. Lemma 2 is a special case of Theorem 6.1 in [65], which is valid not only for invertible states but also non-invertible states. The details of the proof of Lemma 2 and its relation to Theorem 6.1 in [65] can be found in [66].
@@ V QGT determines asymptotic conversion rate
Lemma 6 implies that $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$ if and only if there exists a sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ which asymptotically map $\mathcal{U}_{g}(\psi)^{\otimes N}$ to $\mathcal{U}^{\prime}_{g}(\phi)^{\otimes\lfloor rN\rfloor}$ with vanishing error in $N\to\infty$ for any $g\in G$. Importantly, if we regard $g\in G$ as an unknown parameter, this sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ achieve a conversion of a statistical model $\{\mathcal{U}_{g}(\psi)^{\otimes N}\}_{g\in G}$ to another statistical model $\{\mathcal{U}^{\prime}_{g}(\phi)^{\otimes\lfloor rN\rfloor}\}_{g\in G}$ in the asymptotic limit. The asymptotic behavior of i.i.d. copies of parametric families of states has been studied in the context of QLAN [54, 55, 56, 57], which aims to approximate i.i.d. statistical models by a limit model, called the Gaussian shift model. In the following, after briefly reviewing QLAN for i.i.d. pure-state models in [56, 57], we point out that QGT is a crucial characteristic of their asymptotics. Based on this observation, we construct quantum channels asymptotically converting a pure state to another pure state in RTA, proving Proposition 5.
@@ V.3 Proof of achievability of the conversion rate
The process of transforming $\mathcal{U}_{g}(\psi)^{\otimes N}$ to $\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN\rfloor}$ with asymptotically vanishing error consists of two steps, which we may call the estimation and conversion steps. In the estimation step, we obtain a rough estimate $\hat{g}$ of the true value $g$ by consuming copies of $\mathcal{U}_{g}(\psi)$ sublinear in $N$. In the conversion step, based on the estimate $\hat{g}$, we convert the remaining copies of $\mathcal{U}_{g}(\psi)$ to i.i.d. copies of $\mathcal{U}_{g}^{\prime}(\phi)$ by using the channel described in Lemma 7.
@@
$G_{\mathrm{succ.}}(\delta)\coloneqq\{\hat{g}\in G\mid\exists\bm{\theta},\,\mathcal{U}_{\bm{\theta}}\circ\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi),\,\|\bm{\theta}\|<\delta\}.$ (33)
We define the success probability by
$\mathrm{Pr}_{\mathcal{U}_{g}(\psi)^{\otimes n}}(\|\bm{\theta}\|<\delta)$
$\coloneqq\int_{G}\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes n})\chi_{G_{\mathrm{succ.}}(\delta)}(\hat{g}),$ (34)
@@
We show the following lemma, stating that the existence of an estimator that is successful within an acceptable error of $\delta=N^{-1/2+\epsilon}$ whose failure probability asymptotically vanishes:
@@
Let $G$ be a compact Lie group and $\rho$ be an arbitrary state. Fix $\epsilon\in(0,1/2)$. Then there exists an estimator of $g\in G$ by using consuming $\mathcal{U}_{g}(\rho)^{\otimes n}$ with $n=N^{1-\epsilon}$ such that its success probability satisfies
$\lim_{N\to\infty}\mathrm{Pr}_{\mathcal{U}_{g}(\rho)^{\otimes N^{1-\epsilon}}}(\|\bm{\theta}\|<N^{-1/2+\epsilon})=1.$ (35)
@@
Proposition 5 is proven by combining Lemmas 7 and 8.
@@
From the assumption, $r>0$ satisfies $\mathcal{Q}^{\mathcal{U}_{g}(\psi)}\geq r\mathcal{Q}^{\mathcal{U}_{g}^{\prime}(\phi)}$ for all $g\in G$. Let $\hat{g}\in G$ be the estimator ensured to exist in Lemma 8 for $\epsilon\in(0,1/9)$. Given the estimate $\hat{g}$, we apply Lemma 7 to pure states $(\mathcal{U}_{\hat{g}}(\psi))^{\otimes N^{\prime}}$ and $(\mathcal{U}^{\prime}_{\hat{g}}(\phi))^{\otimes\lfloor rN^{\prime}\rfloor}$ for $N^{\prime}\coloneqq N-N^{1-\epsilon}$ and denote by $\{\mathcal{E}_{N^{\prime}}^{\hat{g}}\}_{N^{\prime}}$ the sequence of conversion channels. Applying $\mathcal{E}_{N^{\prime}}^{\hat{g}}$ after the estimation step, the resulting state is given by
@@
$\coloneqq\int\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes N^{1-\epsilon}})\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right).$ (36)
From the assumption $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$, if parameters $\bm{\theta}$ satisfy $\mathcal{U}_{\bm{\theta}}\circ\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi)$ for some $g,\hat{g}\in G$, then it also holds $\mathcal{U}^{\prime}_{\bm{\theta}}\circ\mathcal{U}^{\prime}_{\hat{g}}(\phi)=\mathcal{U}_{g}^{\prime}(\phi)$. Therefore, for $\hat{g}\in G_{\mathrm{succ.}}(\delta)$ with $\delta>0$, we have
@@
$\coloneqq\inf_{\hat{g}\in G_{\mathrm{succ.}}(N^{-1/2+\epsilon})}\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$\geq\inf_{\hat{g}\in G_{\mathrm{succ.}}((N^{\prime})^{-1/2+\epsilon})}\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$=\inf_{\|\bm{u}\|<(N^{\prime})^{\epsilon}}\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N^{\prime}}}}(\psi_{\hat{g}})^{\otimes N^{\prime}}\right),\mathcal{U}_{\frac{\bm{u}}{\sqrt{N^{\prime}}}}^{\prime}(\phi_{\hat{g}})^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$\xrightarrow{N\to\infty}1,$ (38)
where in the last line, we have used Lemma 7 and the fact that the convergence in trace distance is equivalent to that in fidelity, shown by the Fuchs–van de Graaf inequalities.
@@
$\mathrm{Fid}\left(\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N}),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$=\int_{G}\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes N^{1-\epsilon}})$
@@
$\geq\int\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes N^{1-\epsilon}})\chi_{G_{\mathrm{succ.}}(N^{-1/2+\epsilon})}(\hat{g})$
@@
$\geq\mathrm{Pr}_{\mathcal{U}_{g}(\psi)^{\otimes N^{1-\epsilon}}}(\|\bm{\theta}\|<N^{-1/2+\epsilon})\times f_{N}$
@@
$\lim_{N\to\infty}T\left(\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N}),\mathcal{U}^{\prime}_{g}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)=0$ (40)
holds for any $g\in G$.
@@
$\lim_{N\to\infty}T\left(\Lambda_{N}\circ\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N}),\mathcal{U}^{\prime}_{g}(\phi)^{\otimes\lfloor(r-\delta)N\rfloor}\right)=0$ (41)

### v2

@@ I Introduction
Our formula is valid for any continuous symmetry described by a compact Lie group. To illustrate its wide applicability, we provide four examples. First, since the asymptotic conversion theory among pure states has been completed in prior research exclusively for the case of $U(1)$ [21], we confirm that this result immediately follows from our formula. Second, we provide a necessary and sufficient condition for conversion to be reversible. As a consequence, we affirmatively prove the Marvian-Spekkens conjecture, which had been unsolved for the past decade until today. Third, we analyze examples for $U(1)$ and $SU(2)$ in which the asymptotic conversion is irreversible. The conversion rate for the example in $SU(2)$ is shown to be consistent with the prior study on a specific class of pure states in [2]. Fourth, we calculate the conversion rate for finite groups. Although a finite group is not a Lie group  [^1] Formally, a finite group can be regarded as a Lie group since it can be viewed as a zero-dimensional smooth manifold. However, in our terminology, we consider Lie groups to have a dimension greater than zero., it is shown that our formula is applicable by appending a trivial $U(1)$ phase. We find a divergence in conversion rate, reproducing the prior result on the approximate conversion rate in RTA for finite groups in [34]. Based on our proof of the direct part, we also provide an intuitive explanation of the reason why the approximate conversion rate can diverge in RTA for finite groups.
Our analysis reveals that, for the conversion rate to be nonzero, there must be an inclusion relationship between the symmetry subgroups of the initial and final states. Here, a symmetry subgroup is defined as the set of group transformations that leave a state invariant, serving as a generalization of the period of a state in the $U(1)$ case. However, we show that this stringent constraint imposed by these symmetry subgroups can be circumvented in an alternative scenario where a sublinear number of additional states is available as an auxiliary resource.
Our main result for the conversion rate fully establishes an i.i.d. conversion theory for pure states. We also attempt to extend the analysis to mixed states. Specifically, we study distillable asymmetry and asymmetry cost, which are defined as the optimal rates for converting a mixed state to and from a given pure reference state, respectively. Central to our analysis of distillable asymmetry is a newly introduced asymmetry measure that extends QGT to mixed states. By using the monotonicity of this measure, we derive an upper bound on distillable asymmetry, which generalizes the result in an earlier study [19] in the $U(1)$ case. We also obtain an upper bound on the asymmetry cost and demonstrate that a known result [21] for the $U(1)$ case cannot be directly generalized to arbitrary compact Lie groups.
@@
The rest of this paper is organized as follows: We first review the framework of RTA and the definition of the QGT in Sec. II. The main result of this paper, i.e., the approximate asymptotic conversion rate among pure states in RTA for an arbitrary Lie group, is presented in Sec. III. We also apply our formula for RTA for several setups, providing a unified understanding of prior studies on conversion rates in RTA for different groups [2, 4, 6, 37, 21, 34], and proves the Marvian-Spekkens conjecture [6, 4]. In Sec. IV, we show that the QGT is asymmetry monotone for pure states in one-shot and asymptotic conversion. As a consequence, we prove an inequality providing a bound for the asymptotic conversion rate, i.e., the converse part of the main theorem. In Sec. V, we show the optimality of this bound, i.e., the direct part of the main theorem. In Sec. VI, we prove that a sublinear number of additional states can improve the conversion rate by eliminating the restriction due to symmetry subgroups. In Sec. VII, we investigate mixed-state asymmetry and prove upper bounds on distillable asymmetry and asymmetry cost.
II Preliminaries
We review the resource theory of asymmetry (RTA) and the quantum geometric tensor (QGT), and introduce notations used in this paper.
II.1 Resource theory of asymmetry
@@
II.2 Quantum geometric tensor
@@
for $\mu=1,\cdots,\dim G$, which corresponds to $L(A_{\mu})$, where $L$ is the Lie algebra representation defined by $L(A)\coloneqq-\mathrm{i}\left.\frac{\mathrm{d}}{\mathrm{d}t}U(e^{\mathrm{i}tA})\right|_{t=0}$. Then the QGT for |$\ket{\xi(\bm{\lambda})}$ at |$\ket{\psi}$, which we simply denote by $\mathcal{Q}^{\psi}$, is equal to the non-symmetrized covariance matrix for $\bm{X}\coloneqq\{X_{\mu}\}_{\mu=1}^{\dim G}$:
@@ III Main results
Figure 1: Schematic picture of the setup of Theorem 1, where i.i.d. copies of a pure state $\psi$ are converted into i.i.d. copies of another pure state $\phi$ with an error that vanishes asymptotically.
@@
III.1 U(1) group
$R(\psi\to\phi)=\frac{V(\psi,H)}{V(\phi,H^{\prime})}.$ (9)
III.2 Reversible asymptotic conversion: Proof of the Marvian-Spekkens conjecture
Corollary 2.
  1. (A)
    $\mathrm{Sym}_{G}(\psi)=\mathrm{Sym}_{G}(\phi)$.
  2. (B)
    There uniquely exists $r>0$ such that $\mathcal{Q}^{\psi}=r\mathcal{Q}^{\phi}$.
Proof.
[…]
  1. (i)
    $\mathrm{Sym}_{G}(\psi)=\mathrm{Sym}_{G}(\phi)$.
  2. (ii)
    $C_{\mathfrak{g}}(\psi)=R(\psi\to\phi)C_{\mathfrak{g}}(\phi)$.
  3. (iii)
[…]
III.3 Irreversibility in asymptotic conversion
$\mathcal{M}(\psi)$ $\coloneqq 2\braket{\psi}{\mathcal{J}}{\psi},$ (14)
[…]
Since the QGT is calculated for $\psi$ as
III.4 Finite groups
reproducing the result on the approximate asymptotic conversion rate in [34]. We remark that our construction of the conversion channels provides an intuitive explanation of the reason why the conversion rate for finite groups diverges, as will be explained at the end in Sec. V.3.
@@ IV.1 Monotonicity of QGT in one-shot conversion
$\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}=\frac{1}{1-q}V(\psi,O)+\frac{1}{q}V(\psi,O^{\dagger}),$ (21)
@@
In literature, the proof of the monotonicity of monotone metric is sometimes provided under the assumption that $\rho$ is invertible. Lemma 3 is a special case of Theorem 6.1 in [68], which is valid not only for invertible states but also non-invertible states. The details of the proof of Lemma 3 and its relation to Theorem 6.1 in [68] are presented in Appendix B.1.
@@ V QGT determines asymptotic conversion rate
Lemma 7 implies that $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$ if and only if such that
$\lim_{N\to\infty}\sup_{g\in G}T\left(\mathcal{E}_{N}\left(\mathcal{U}_{g}(\psi)^{\otimes N}\right),\mathcal{U}^{\prime}_{g}(\phi)^{\otimes\lfloor rN\rfloor}\right)=0.$ (39)
Importantly, if we regard $g\in G$ as an unknown parameter, this sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ achieve a conversion of a statistical model $\{\mathcal{U}_{g}(\psi)^{\otimes N}\}_{g\in G}$ to another statistical model $\{\mathcal{U}^{\prime}_{g}(\phi)^{\otimes\lfloor rN\rfloor}\}_{g\in G}$ in the asymptotic limit. The asymptotic behavior of i.i.d. copies of parametric families of states has been studied in the context of QLAN [54, 55, 56, 57], which aims to approximate i.i.d. statistical models by a limit model, called the Gaussian shift model. In the following, after briefly reviewing QLAN for i.i.d. pure-state models in [56, 57], we point out that QGT is a crucial characteristic of their asymptotics. Based on this observation, we construct quantum channels asymptotically converting a pure state to another pure state in RTA, proving Proposition 6.
@@ V.3 Proof of achievability of the conversion rate
The process of transforming $\mathcal{U}_{g}(\psi)^{\otimes N}$ to $\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN\rfloor}$ with asymptotically vanishing error consists of two steps, which we may call the estimation and conversion steps. In the estimation step, we obtain a rough estimate $\hat{g}$ of the true value $g$ by consuming copies of $\mathcal{U}_{g}(\psi)$ sublinear in $N$. In the conversion step, based on the estimate $\hat{g}$, we convert the remaining copies of $\mathcal{U}_{g}(\psi)$ to i.i.d. copies of $\mathcal{U}_{g}^{\prime}(\phi)$ with quantum channels constructed by using Lemma 8.
@@
$G_{\mathrm{succ.}}^{(g,\delta)}\coloneqq\{\hat{g}\in G\mid\exists\bm{\theta},\,\mathcal{U}_{\bm{\theta}}\circ\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi),\,\|\bm{\theta}\|<\delta\}.$ (43)
We quantify the success probability in the worst case by
$p^{\mathrm{succ.}}(\delta)\coloneqq$ $\inf_{g\in G}\int_{G}\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes n})\chi_{G_{\mathrm{succ.}}^{(g,\delta)}}(\hat{g}),$ (44)
@@
We show the following lemma, which guarantees the existence of an estimator that is successful within an acceptable error of $\delta=N^{-1/2+\epsilon}$ whose failure probability asymptotically vanishes:
@@
Let $G$ be a compact Lie group and $\rho$ be an arbitrary state. Fix $\epsilon\in(0,1/2)$. Then there exists an estimator of $g\in G$, which consumes $\mathcal{U}_{g}(\rho)^{\otimes n}$ with $n=\lceil N^{1-\epsilon}\rceil$, such that its worst-case success probability satisfies
$\lim_{N\to\infty}p^{\mathrm{succ.}}(N^{-1/2+\epsilon})=1.$ (45)
@@
In the conversion step, we convert states using quantum channels guaranteed by the following lemma:
Lemma 10.
Let $U,U^{\prime}$ be (non-projective) unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Let |$\ket{\psi}\in\mathcal{H}$ and |ϕketitalic-$\ket{\phi}\in\mathcal{H}^{\prime}$ be pure states. We define $\psi_{g}\coloneqq\mathcal{U}_{g}(\psi)$ and $\phi_{g}\coloneqq\mathcal{U}^{\prime}_{g}(\phi)$ and denote their QGTs by $\mathcal{Q}^{\psi_{g}}$ and $\mathcal{Q}^{\phi_{g}}$. If $r>0$ satisfies $\mathcal{Q}^{\psi_{g}}\geq r\mathcal{Q}^{\phi_{g}}$ for all $g\in G$, then, there exists a sequence of quantum channels $\{\mathcal{E}^{(g)}_{N}\}_{N}$ such that the conversion error
$\delta_{N}(g,\bm{u})\coloneqq T\left(\mathcal{E}^{(g)}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\left(\psi_{g}\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}\left(\phi_{g}\right)^{\otimes\lfloor rN\rfloor}\right)$ (46)
satisfies $\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon}}\delta_{N}(g,\bm{u})=0$ for $\epsilon\in(0,1/9)$.
This lemma is proven from Lemma 8 by using properties of a compact Lie group. The proof is detailed in Appendix D.
Proposition 6 is proven by combining Lemmas 9 and 10.
@@
From the assumption, $r>0$ satisfies $\mathcal{Q}^{\mathcal{U}_{g}(\psi)}\geq r\mathcal{Q}^{\mathcal{U}_{g}^{\prime}(\phi)}$ for all $g\in G$. Let $\hat{g}\in G$ be the estimator ensured to exist in Lemma 9 for $\epsilon\in(0,1/9)$. Given the estimate $\hat{g}$, we apply Lemma 10 to pure states $(\mathcal{U}_{\hat{g}}(\psi))^{\otimes N^{\prime}}$ and $(\mathcal{U}^{\prime}_{\hat{g}}(\phi))^{\otimes\lfloor rN^{\prime}\rfloor}$ for $N^{\prime}\coloneqq N-\lceil N^{1-\epsilon}\rceil$ and denote by $\{\mathcal{E}_{N^{\prime}}^{(\hat{g})}\}_{N^{\prime}}$ the sequence of conversion channels. Applying $\mathcal{E}_{N^{\prime}}^{(\hat{g})}$ after the estimation step, the resulting state is given by
@@
$\coloneqq\int\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes\lceil N^{1-\epsilon}\rceil})\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right).$ (47)
From the assumption $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$, if parameters $\bm{\theta}$ satisfy $\mathcal{U}_{\bm{\theta}}\circ\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi)$ for some $g,\hat{g}\in G$, then it also holds $\mathcal{U}^{\prime}_{\bm{\theta}}\circ\mathcal{U}^{\prime}_{\hat{g}}(\phi)=\mathcal{U}_{g}^{\prime}(\phi)$. Therefore, for $\hat{g}\in G_{\mathrm{succ.}}^{(g,\delta)}$ with $\delta>0$, we have
@@
$\coloneqq\inf_{g\in G}\inf_{\hat{g}\in G_{\mathrm{succ.}}^{(g,N^{-1/2+\epsilon})}}\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$\geq\inf_{g\in G}\inf_{\hat{g}\in G_{\mathrm{succ.}}^{(g,(N^{\prime})^{-1/2+\epsilon})}}\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$=\inf_{\hat{g}\in G}\inf_{\|\bm{u}\|<(N^{\prime})^{\epsilon}}\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N^{\prime}}}}(\psi_{\hat{g}})^{\otimes N^{\prime}}\right),\mathcal{U}_{\frac{\bm{u}}{\sqrt{N^{\prime}}}}^{\prime}(\phi_{\hat{g}})^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$\geq\left(1-\sup_{\hat{g}\in G}\sup_{\|\bm{u}\|<(N^{\prime})^{\epsilon}}\delta_{N^{\prime}}(\hat{g},\bm{u})\right)^{2}\xrightarrow{N\to\infty}1,$ (49)
where in the last line, we used the Fuchs–van de Graaf inequalities to relate the fidelity to the trance distance $\delta_{N}(g,\bm{u})$ defined in Eq. (46) in Lemma 10.
@@
$\inf_{g\in G}\mathrm{Fid}\left(\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N}),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$=\inf_{g\in G}\int_{G}\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes\lceil N^{1-\epsilon}\rceil})$
@@
$\geq\inf_{g\in G}\int\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes\lceil N^{1-\epsilon}\rceil})\chi_{G_{\mathrm{succ.}}^{(g,N^{-1/2+\epsilon})}}(\hat{g})$
@@
$\geq p^{\mathrm{succ.}}(N^{-1/2+\epsilon})\times f_{N}$
@@
$\lim_{N\to\infty}\sup_{g\in G}T\left(\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N}),\mathcal{U}^{\prime}_{g}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)=0.$ (51)
@@
$\lim_{N\to\infty}\sup_{g\in G}T\left(\Lambda_{N}\circ\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N}),\mathcal{U}^{\prime}_{g}(\phi)^{\otimes\lfloor(r-\delta)N\rfloor}\right)=0.$ (52)